### v1

# Prospect Theory Fails for LLMs: Revealing Instability of Decision-Making under Epistemic Uncertainty

Abstract
Prospect Theory (PT) models human decision-making under uncertainty, while epistemic markers (e.g., “maybe”) serve to express uncertainty in language. However, it remains largely unexplored whether Prospect Theory applies to contemporary Large Language Models and whether epistemic markers, which express human uncertainty, affect their decision-making behavior. To address these research gaps, we design a three-stage experiment based on economic questionnaires. We propose a more general and precise evaluation framework to model LLMs’ decision-making behavior under PT, introducing uncertainty through the empirical probability values associated with commonly used epistemic markers in comparable contexts. We then incorporate epistemic markers into the evaluation framework based on their corresponding probability values to examine their influence on LLM decision-making behaviors. Our findings suggest that modeling LLMs’ decision-making with PT is not consistently reliable, particularly when uncertainty is expressed in diverse linguistic forms. Our code is released in https://github.com/HKUST-KnowComp/MarPT.
Prospect Theory Fails for LLMs: Revealing Instability of Decision-Making under Epistemic Uncertainty
Rui Wangthanks: Equal contribution.    Qihan Lin    Jiayu Liu    Qing Zong    Tianshi Zheng    Weiqi Wang    Yangqiu Song Hong Kong University of Science and Technology Huazhong University of Science and Technology rwangcn@connect.ust.hk       yqsong@cse.ust.hk
Figure 1: An overview of our three-stage experiment. Stage 1 fits PT parameters from binary choices with precise probabilities. Stage 2 estimates each marker’s internal probability from the point where options K and U are chosen equally. Stage 3 substitutes probabilities with markers to measure the effect of linguistic uncertainty.
1 Introduction
Human decision-making under uncertainty has been extensively studied through both normative and descriptive frameworks, among which Prospect Theory (PT) Kahneman and Tversky (1979); Tanaka et al. (2010); Rathi et al. (2025) stands out as particularly influential. PT models a range of behavioral patterns by introducing psychologically grounded parameters, such as loss aversion and probability weighting P. and H. (2005); Gneezy et al. (2006); Rathi et al. (2025). Given the increasing deployment of LLMs in uncertainty-sensitive applications Wei et al. (2025, 2024); Phan et al. (2025), it is crucial to examine whether their decision behaviors align with PT, as such alignment has been shown to enhance human-AI collaboration efficiency Willems et al. (2019); Yao et al. (2024).
In parallel, human uncertainty expressions are commonly expressed via epistemic markers to support flexible language use Belem et al. (2024); Liu et al. (2025b); Zhou et al. (2023). Despite their ubiquity, these expressions are inherently ambiguous Agata (2017), and correctly interpreting them is essential for effective decision-making Mahmoodi et al. (2020). Traditional decision-making frameworks Laplace (1812); Bernoulli (1738); Von Neumann and Morgenstern (1944); Roy (1968) often fail to account for this form of linguistic uncertainty, resulting in a mismatch between computational models and real-world reasoning. While humans are capable of intuitively navigating such expressions Wallsten (1986); Erev and Cohen (1990); Kadavath et al. (2022), it remains unclear how LLMs interpret and respond to them in decision-making contexts.
Existing studies Jia et al. (2024); Liu et al. (2025a, 2024) often assume that LLMs’ decision-making can be directly modeled using PT, without systematically evaluating the suitability of this human-centered psychological framework. Moreover, it remains unexplored whether LLMs exhibit consistent decision-making behavior when probabilities are conveyed through epistemic markers, as is common in real-world communication.
To examine how LLMs handle decision-making under uncertainty, we design a three-stage lottery-based evaluation adapted from classic behavioral economics Allais (1953); Brandstätter et al. (2006); Tanaka et al. (2010). Our evaluation consists of three stages, illustrated in Figure 1. We begin by presenting binary choice tasks with precise numerical probabilities to estimate each model’s risk preferences and fit the PT parameters. In the second stage, we introduce linguistic uncertainty by replacing the numerical probabilities with epistemic markers, allowing us to infer the model’s implicit probabilistic interpretation of these expressions. Finally, we reassess model behavior on the original decision tasks, now framed using epistemic markers grounded in their inferred probability values, to directly evaluate the impact of linguistic uncertainty on decision-making.
Our results reveal three key insights. First, smaller models fail to capture PT-like behavior, whereas larger models (e.g., Qwen2.5-32B-Instruct) demonstrate better alignment with PT predictions. Second, models vary substantially in their interpretations of epistemic markers, resulting in divergent probability mappings. Third, introducing epistemic markers disrupts decision consistency and alters PT parameters, exposing the fragility of LLMs’ decision-making under linguistic uncertainty. These findings point to fundamental gaps in the interpretability and robustness of current LLMs when processing ambiguous language, and underscore the need for improved epistemic calibration and more nuanced theoretical frameworks for modeling LLM behavior in uncertain contexts.
2 Related Work
Prospect Theory in Economic Decision-Making.
Prospect Theory Kahneman and Tversky (1979) has long served as a foundational framework for modeling human decision-making under risk, capturing key behavioral patterns like loss aversion and probability distortion. Empirical studies extend PT to diverse populations and settings Tanaka et al. (2010), and recent work explores whether LLMs exhibit similar patterns Jia et al. (2024). Our work builds on these efforts but focuses on how language-based uncertainty impacts PT-consistent behavior in LLMs.
LLM Decision-Making under Uncertainty.
Recent work evaluates how LLMs handle uncertain scenarios, including economic games, moral dilemmas, and ambiguous instructions Liu et al. (2025a); Jia et al. (2024); Zong et al. (2025); Zheng et al. (2025d). These studies often probe whether LLMs mimic human cognitive biases or align with normative models. We extend this line of inquiry by testing whether linguistic uncertainty conveyed by epistemic markers causes instability of LLM behaviors in decision-making situations.
Epistemic Markers in NLP.
Recent work investigates how LLMs interpret and respond to linguistic signals related to uncertainty and confidence. Zhou et al. (2023) study how different epistemic markers embedded in prompts affect model predictions. Belem et al. (2024) evaluate LLMs’ interpretation of epistemic markers, finding partial human-like behavior but systematic biases. Liu et al. (2025b) further argue that such markers are often unreliable indicators of internal confidence in LLMs. These findings suggest that models may mimic surface linguistic patterns rather than demonstrate true epistemic reasoning. Our study builds on this line of work by examining how LLMs understand epistemic markers in economic contexts with uncertainty.
[…]
Figure 2: Visual illustration of the three PT parameters ($\sigma$, $\lambda$, $\gamma$). Each parameter is shown with its meaning and an interpretation of its directional significance.
3 Preliminary
In rational decision theory, an agent’s preference over risky prospects follows the von Neumann-Morgenstern expected utility framework. For a prospect $P=(x_{1},p_{1};\cdots;x_{n},p_{n})$ yielding outcome $x_{i}$ with probability $p_{i}$, the expected utility is computed as:
$EU(P)=\sum_{i=1}^{n}p_{i}\cdot u(x_{i}),$ (1)
where $u(\cdot)$ is a cardinal utility function mapping outcomes to real numbers. Human decisions are assumed to maximize $EU(P)$ under traditional Expected Utility Theory (EUT), given by von Neumann et al. (1944). However, empirical evidence systematically violates EUT assumptions. Prospect Theory (PT) addresses these anomalies through three psychological distortions of rational utility. It maintains a utility-based approach but fundamentally alters Expected Utility Theory through three key properties:
  • Risk Preference ($\sigma$): Agents often exhibit varying degrees of risk aversion or risk-seeking behavior.
  • Loss Aversion ($\lambda$): Losses psychologically outweigh equivalent gains.
  • probability weighting ($\gamma$): Agents often exhibit systematic probability distortion, overweighting small probabilities and underweighting high probabilities.
To capture these characteristics, Prospect Theory introduce two specialized functions: the value function formalizes how the outcomes translate into subjective utility, while the probability weighting function captures non-linear probability perception. Together, these functions model the distorted utility calculations that characterize PT decision-making.
The value function v(x) quantifies subjective satisfaction from outcomes relative to a reference point (zero in this study). For outcomes $x\geq 0$ (gains) and $x<0$ (losses), we adopt the parameterized form:
$v(x)=\begin{cases}x^{\sigma}&\text{for }x\geq 0\\ -\lambda(-x)^{\sigma}&\text{for }x<0.\end{cases}$ (2)
Key parameters here are loss aversion ($\lambda$) which is a multiplier of negative perceptions, and risk preference ($\sigma$) which controls curvature (sensitivity to values).
The probability weighting function formalizes the transformation of objective probabilities $p$ in subjective decision weights. We adopt the Kahneman and Tversky (1979) probability weighting function:
$w(p)=\frac{p^{\gamma}}{\left(p^{\gamma}+(1-p)^{\gamma}\right)^{1/\gamma}},$ (3)
where $\gamma$ controls the curvature of the function.
The final utility for any binary prospect in the form $P=(x,p;y,q)$ is defined as follows:
$u(P)=\begin{cases}v(y)+w(p)(v(x)-v(y))&\mkern-15.0mu[1]\\ w(p)v(x)+w(q)v(y)&\mkern-15.0mu[2],\end{cases}$ (4)
where $[1]$ is $\text{if }x>y>0\text{ or }x<y<0$” and $[2]$ is $\text{if }x<0<y$”.
4 Decision-Making Behavior Evaluation
Based on the above economic frameworks, we mainly adopt the three-series lottery-choice experiment developed by Tanaka et al. (2010) to get a reliable PT parameter measurement. Series 1 and 2 are designed to elicit risk preference ($\sigma$) and probability weighting ($\gamma$), while series 3 is designed for loss aversion ($\lambda$). The prospect settings are shown in Appendix A, and the prompt design is in Appendix B.
Each lottery consists of two option: a relatively safer option K with lower risk, and a riskier option U. The agent is asked to directly choose from option K and U based on its risk preference.
After sampling 256 times for each question, we count the portion of choosing option K for each lottery. Then we define the predicted probability of choosing option K for each lottery as follows:
[…]
where $\text{EUDIFF}=u(\text{K})-u(\text{U})$ is the difference in the distorted utility of prospect K and U under prospect theory. We choose the sigmoid function as it is a usual setting in economic studies, e.g., Chakravarty and Roy (2009). We add up the Bernoulli log-likelihood for all 35 lotteries as the negative log-likelihood function, and run MLE with this function to estimate $\sigma,\lambda\text{ and }\gamma$.
To get confidence intervals, we use a bootstrap method Efron (1979) by generating simulated datasets through binomial sampling from predicted probabilities derived from the original parameter estimates. Specifically, for each observation $i$, we sample $\tilde{y}_{i}\sim\text{Binomial}(n=1,p=\hat{p}_{i})$ where $\hat{p}_{i}$ is the predicted probability. The model parameter standard deviation $\sigma_{\hat{\theta}}$ is estimated from the bootstrap distribution, and the 95% confidence interval is constructed using the percentile method:
$\text{CI}_{95\%}=\left[\hat{\theta}_{(0.025)}^{*},\hat{\theta}_{(0.975)}^{*}\right],$ (6)
where $\hat{\theta}_{(\alpha)}^{*}$ denotes the $\alpha$-quantile of the bootstrap parameter estimates. This approach accounts for parameter uncertainty in finite samples.
We then quantify model goodness-of-fit by computing McFadden pseudo-$\mathbf{R^{2}}$ McFadden (1977), defined as:
$R^{2}_{\text{McFadden}}=1-\frac{\mathcal{L}_{\text{PT}}}{\mathcal{L}_{\text{null}}},$ (7)
where $\mathcal{L}_{\text{PT}}$ is the log-likelihood of Prospect Theory and $\mathcal{L}_{\text{null}}$ represents the log-likelihood of the intercept-only model with uniform choice probabilities. This metric measures the improvement of our model over random guessing.
Finally, we calculate the mean absolute error (MAE) between the actual probability $p_{\text{actual}}$ and the predicted probability $p_{\text{pred}}$ of choosing option K derived from our PT model:
$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}\left|p_{\text{actual}}^{(i)}-p_{\text{pred}}^{(i)}\right|,$ (8)
where $N$ denotes the number of observations. This provides a direct measure of prediction error.
Overall, the three PT parameters, their confidence intervals, the McFadden pseudo-$R^{2}$, and the MAE score together describe the revealed agent’s risk attitude and our model’s explanatory power itself.
5 Decision-making Scenarios with Uncertainty
One of the most fundamental yet underexplored challenges in aligning AI with human reasoning lies in how uncertainty is communicated and interpreted. This section studies decision-making when verbal probability expressions—epistemic markers—replace numerical probabilities. Section 5.1 estimates their numerical equivalents via a controlled lottery experiment, and Section 5.2 applies these values in the PT framework to re-measure PT parameters. The prompt design is in Appendix B.
5.1 Probability Mapping of Epistemic Markers
Epistemic markers are inherently vague and context-sensitive, yet they often substitute for precise numerical probabilities in practice. For LLMs, the ability to interpret epistemic markers consistently and meaningfully is critical if they are used as decision-support tools. However, there is little empirical understanding of how LLMs map epistemic markers to numerical probabilities, and whether this mapping is coherent across different models or aligned with human intuition.
No. Epistemic Marker
Probability Mapping
by Human
1 almost certain 95%
2 highly likely 90%
3 very likely 90%
4 likely 80%
5 probable 70%
6 somewhat likely 70%
7 possible 60%
8 uncertain 50%
9 somewhat unlikely 30%
10 unlikely 25%
11 not likely 20%
12 doubtful 20%
13 very unlikely 10%
14 highly unlikely 10%
Table 1: Epistemic markers used in the experiment. Human probability mapping comes from Belem et al. (2024)
To address this issue, we design a controlled lottery experiment in an economic decision-making context. Each trial presents the model with a choice between two options. For option K, there is a fixed probability pp% of winning \$100, while $p$ traverses all values in $probs=[5,15,25,35,45,55,65,75,85,95]$. For option U, there is an unknown possibility of winning \$100 which is defined by the different markers. We select 14 commonly used markers from previous work in Belem et al. (2024), and present them in Table 1.
For each marker, we record the number of times the model selects option K, denoted as $\text{NUM}K$. The key idea is: when $\text{NUM}K$ reaches 50% of the total trials for a given probability $p$, the model considers the two options equally attractive. We define the mapping probability $p_{mapping}$ for that marker as the value of $p$ at which this occurs:
$p_{mapping}=p.$ (9)
Figure 3: An illustration of calculating $p_{mapping}$. We use linear interpolation for fitting and calculate $p_{mapping}$ through the similarity ratio.
Since $p$ is sampled at discrete points, the exact 50% point may fall between two sampled probabilities. To estimate $p_{\text{mapping}}$, we perform linear interpolation between the two nearest points. Let $n_{0}=0.5\cdot sample\_number$ be the target count corresponding to 50% selection of option K, $(p_{x},cnt_{x})$ be the probability-count pair just below $n_{0}$, and $(p_{y},cnt_{y})$ be the pair just above $n_{0}$. As illustrated in Figure 3, by the slope formula
$\frac{p_{mapping}-p_{x}}{p_{y}-p_{x}}=\frac{n_{0}-cnt_{x}}{cnt_{y}-cnt_{x}},$ (10)
we solve for $p_{mapping}$:
$p_{mapping}=\frac{(n_{0}-cnt_{x})\cdot p_{y}+(cnt_{y}-n_{0})\cdot p_{x}}{cnt_{y}-cnt_{x}}.$ (11)
Through this experiment, we obtain a list of 14 probability values (one per marker) for each model, capturing how it semantically interprets verbal uncertainty in economic terms.
5.2 Re-measurement of Prospect Theory Parameters
After we have established the probability mapping, we pick a pair of epistemic markers which has the closest normalized probability to the original settings. This is to ensure that the probability setting does not change much, so that the condition is controlled. For example, for a model with “somewhat unlikely” mapped to $p_{1}=32\%$ and “highly likely” mapped to $p_{2}=68\%$, we normalize them using the formula:
$p^{\prime}_{1}=\frac{p_{1}}{p_{1}+p_{2}},\quad p^{\prime}_{2}=\frac{p_{2}}{p_{1}+p_{2}}$ (12)
to ensure they add up to $100\%$. Then we replace the probabilities in the original decision-making behavior evaluation framework with the closest epistemic marker pairs. We re-run the PT metrics measurement test using $p^{\prime}_{1}$ and $p^{\prime}_{2}$ and compare the result with the original study. There are four rounds in our study. We control the option to which epistemic markers are introduced. In round 1, we only introduce markers in option K, series 1 and 2. In round 2, all 3 series of option K are with epistemic markers. In round 3, all option U are with markers. In round 4, both option K and option U of all series are with epistemic markers.
[…]
Human 0.670 2.630 0.685 - -
Llama-3.1-8B-Instruct 0.585 0.010 0.753 0.332 0.092
Mistral-7B-Instruct-v0.3 0.534 0.570 0.577 0.155 0.132
Qwen2.5-7B-Instruct 0.429 0.010 3.645 0.047 0.116
Qwen2.5-14B-Instruct 0.503 1.909 0.896 0.257 0.067
Qwen2.5-32B-Instruct 0.598 1.213 0.867 0.161 0.225
Table 2: Baseline PT parameter estimation across LLMs. Human reference values are taken from Tanaka et al. (2010). Most LLMs exhibit human-like risk aversion and probability distortion, but show inconsistent and generally weaker loss aversion. We bold the results where MAE $\leq 0.20$ and $R^{2}\geq 0.10$, indicating that PT provides a reliable fit.
6 Results and Findings
Model Top 7 Epistemic Markers
almost certain highly likely very likely likely probable somewhat likely possible
Llama-3.1-8B-Instruct 87.92 56.04 58.00 41.80 44.23 36.71 36.29
Mistral-7B-Instruct-v0.3 96.80 67.89 63.10 57.50 87.22 48.98 52.16
Qwen2.5-7B-Instruct 82.71 67.00 67.06 04.78 03.44 08.93 04.51
Qwen2.5-14B-Instruct 91.56 55.00 54.10 42.38 26.51 32.47 38.38
Qwen2.5-32B-Instruct 97.50 95.08 82.82 65.00 54.74 46.08 55.00
Table 3: Switching probabilities (%) for top 7 epistemic markers across models. Models differ substantially in how they interpret high-certainty markers such as almost certain or very likely, with mapping values varying by over 50 percentage points across models, indicating inconsistent internal representations of certainty.
Model Bottom 7 Epistemic Markers
uncertain somewhat unlikely unlikely not likely doubtful very unlikely highly unlikely
Llama-3.1-8B-Instruct 35.49 33.71 33.49 36.91 33.65 31.30 32.83
Mistral-7B-Instruct-v0.3 48.08 40.15 38.27 30.93 34.03 29.47 27.88
Qwen2.5-7B-Instruct 35.58 27.24 19.90 27.70 25.69 18.37 19.32
Qwen2.5-14B-Instruct 29.03 26.45 19.10 20.82 13.04 10.94 10.52
Qwen2.5-32B-Instruct 02.98 21.89 03.33 03.08 03.42 02.77 02.51
Table 4: Switching probabilities (%) for bottom 7 epistemic markers across models. Several models compress distinct low-certainty markers into similarly low probabilities, especially in larger models like Qwen2.5-32B-Instruct, suggesting a limited sensitivity to fine-grained uncertainty.
6.1 Risk Preference under Numerical Probabilities
For decision-making under exact probability, we obtain different result compared with previous works by Jia et al. (2024) and Liu et al. (2025a). The key parameter values we obtained are shown in Table 2, and more detail can be found in Appendix C.
Not all LLMs decision-making choices can be explained by Prospect Theory.
Following human econometric model regression standards, we consider a MAE score >> 0.20 means regression result not reliable, and a McFadden pseudo-$R^{2}$ score << 0.10 means using PT may be meaningless. McFadden (1977) Llama-3.1-8B-Instruct and Qwen2.5-14B-Instruct have too high MAE and too low $R^{2}$, which indicates they do not fit into Prospect Theory well.
LLMs are consistent and show human-like risk preference, but not consistent and show significantly lower loss aversion than human.
All models show some extent of risk aversion $\sigma$ between 0.4 and 0.6 (a bit lower than human), but they are not consistent in $\lambda$ value and all significantly lower than human. Moreover, some models are more sensitive to loss ($\lambda>1$), while others seems more sensitive to gain($\lambda<1$). Llama-3.1-8B-Instruct and Qwen2.5-7B-Instruct have unusual boundary $\lambda$ values, which implies they have an irregular choice pattern.
Majority of LLMs show human-like probability distortion.
We observe that the majority of human-like probability weighting in LLMs, which is overweighting small probabilities and underweighting large probabilities. This is the opposite result given by Jia et al. (2024).
Figure 4: Probability mapping of different models for different markers. While different models assign divergent probability values to the same marker, the relative ordering of markers remains broadly consistent.
6.2 Cross-Model Comparison of Marker Mappings
The experimental results are shown in Tables 3 and 4. We also intuitively present the probability mapping of different models for different markers in Figure 4.
Different models assign different probabilities to the same epistemic markers.
Across the top 7 markers, the mapping probabilities vary significantly between models. For example, almost certain is mapped to over 97% in Qwen2.5-32B-Instruct but only less than 83% in Qwen2.5-7B-Instruct, while somewhat likely ranges from 8.9% (Qwen2.5-7B-Instruct) to 48.9% (Mistral-7B-Instruct-v0.3). This highlights that language models, even within the same family, do not share a consistent internal representation of uncertainty expressions.
Despite the divergence in absolute values, there is a striking consistency in relative ordering among markers.
Nearly all models assign the highest probabilities to “almost certain”, followed by “highly likely” and “very likely”, suggesting that models broadly agree on the ordinal semantics of epistemic language even if they differ numerically. This partial consistency could be useful for comparative or ranking-based tasks, but it is insufficient for applications requiring calibrated probabilistic reasoning.
Some models collapse multiple distinct markers into indistinguishably low probabilities.
Some models exhibit a “compression” effect, where multiple markers are mapped to similar probabilities. For instance, Qwen2.5-32B-Instruct maps 6 weakest markers to probabilities between 2.5% and 3.5%, effectively collapsing fine-grained uncertainty distinctions. Such behavior suggests an overconfident or categorical interpretation of ambiguous language, which undermines the expressiveness of epistemic markers. Similarly, Llama-3.1-8B-Instruct displays a narrow dynamic range for markers, most of them close to 30%, indicating limited sensitivity of epistemic markers.
Larger model size does not guarantee human-aligned probability interpretations.
Interestingly, larger models do not always produce more calibrated mappings. Although Qwen2.5-32B-Instruct maps a wider range, with some close to 100% and some near 0%, this does not prevent it from having poor discrimination on low-probability markers. Meanwhile, smaller models such as Mistral-7B-Instruct-v0.3 perform no worse than Qwen2.5-32B-Instruct across the entire mapping of markers. These patterns suggest that scale alone does not guarantee a nuanced probabilistic understanding of epistemic markers, and that other factors such as model architecture may play a more decisive role.
6.3 Change of Decision-Making Behavior under Epistemic Markers
The estimation results under varying degrees of linguistic uncertainty are illustrated in Figure 5 and 6. We examine how replacing precise numerical probabilities with epistemic markers affects LLMs’ internal decision-making parameters and performance under the framework of Prospect Theory. The analysis covers four experimental conditions, including marker substitution in either or both options of the lottery choice task.
Figure 5: The PT parameters estimated across rounds. The “Baseline” represents results from the first stage of our experiment. In subsequent rounds, parameters fluctuate relative to the baseline, indicating instability in the model’s PT alignment under epistemic perturbations.
Figure 6: The PT performance across rounds. Similar to Table 5, the “Baseline” displays each model’s original PT fit. Across later rounds, both MAE and McFadden $R^{2}$ values vary considerably, further suggesting that LLM behavior under uncertainty is not reliably captured by human cognitive models.
Linguistic uncertainty causes moderate shifts in risk preference but affects other decision parameters more profoundly.
For the majority of models, the estimated risk preference remains relatively stable despite replacing numeric probabilities with epistemic markers, indicating that risk attitudes are only mildly perturbed by uncertainty. However, models differ considerably in their loss aversion and probability weighting under these conditions. Larger models tend to maintain more consistent loss sensitivity resembling human behavior, whereas smaller models show larger fluctuations. Similarly, the increase in $\gamma$ after adding disturbance shows that most models exhibit increasingly conservative probability distortions, reflecting cautious decision patterns under ambiguity.
Epistemic markers sometimes activate a more PT-like decision-making behavior for LLMs.
This may because LLMs are more sensitive to linguistic uncertainty than numeric probability value, showing a different probability interpretation of epistemic markers as we obtain in stage 2. However, this finding does not mean epistemic markers can guarantee a more PT-like decision-making behavior. It further strengthen our claim that LLMs’ decision-making is not stable and explainable.
Larger size LLMs seem to have more stable decision-making behavior under linguistic uncertainty.
While smaller language models (e.g. Mistral-7B-Instruct-v0.3) exhibit drastically varying PT parameters under epistemic markers, sometimes even touch boundary values (e.g. close to zero loss aversion), which indicates a fail regression. Larger models (e.g. Qwen2.5-32B-Instruct) are more robust to linguistic uncertainty.
7 Conclusion
This study challenges the direct application of human-centric Prospect Theory to LLMs, especially in linguistically uncertain contexts. Through a novel three-stage framework combining behavioral economics and semantic analysis, we show that LLMs (1) may not always fit into Prospect Theory (2) exhibit widely divergent interpretations to epistemic markers under decision-making context (3) fail to maintain stable and explainable risk preferences under epistemic markers despite scale advantages. We show that PT fails to model LLMs with epistemic uncertainty and present the first mapping between markers and economic decisions. Future work may explore AI-specific decision theories beyond human dicision models and calibration methods for epistemic markers.
Limitations
We deliberately choose non-reasoning models and ask them to directly make their choice, because nearly all models tend to directly calculate the expected utility and choose the option with higher EU if they are allowed to reason. This contradicts with our purpose, which is to obtain their risk attitude under quick decisions. However, avoiding reasoning models limits the applicability of our results. Meanwhile, we do not find a universal pattern for PT parameter changes under epistemic markers. Another limitation is that the probability mapping for epistemic markers are obtained in a slightly different context with the decision-making evaluation test. Additionally, our tests are all done in financial decision making context, which may limits its versatility.
Acknowledgments
We thank the anonymous reviewers and the area chair for their constructive comments. The authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), the RIF (R6021-20) and the GRF (16205322) from RGC of Hong Kong,SAR, China.
Appendix A Lottery Design
Table 5, 6 and 7 shows the lottery design for PT parameter estimation. The values here are from Tanaka et al. (2010). They are specially designed to get best PT parameters.
Option K Option U
Lottery 30% 70% 10% 90%
1 40 10 68 5
2 40 10 75 5
3 40 10 83 5
4 40 10 93 5
5 40 10 106 5
6 40 10 125 5
7 40 10 150 5
8 40 10 185 5
9 40 10 220 5
10 40 10 300 5
11 40 10 400 5
12 40 10 600 5
13 40 10 1000 5
14 40 10 1700 5
Table 5: Series 1 both options are gains.
Option K Option U
Lottery 90% 10% 70% 30%
1 40 30 54 5
2 40 30 56 5
3 40 30 58 5
4 40 30 60 5
5 40 30 62 5
6 40 30 65 5
7 40 30 68 5
8 40 30 72 5
9 40 30 77 5
10 40 30 83 5
11 40 30 90 5
12 40 30 100 5
13 40 30 110 5
14 40 30 130 5
Table 6: Series 2 both options are gains.
Option K Option U
50% 50% 50% 50%
Lottery Win Lose Win Lose
1 25 4 30 21
2 4 4 30 21
3 1 4 30 21
4 1 4 30 16
5 1 8 30 16
6 1 8 30 14
7 1 8 30 11
Table 7: Series 3 both options have gains and losses.
Appendix B Prompt Design
Figure 7 shows prompt design for decision-making evaluation test and probability mapping test. All lotteries are sampled 256 times. To simulate human decision-making, while keeping the model directly output its answer, up to 15 history decisions are maintained. Meanwhile, we use a random order for the 35 lotteries to relieve positional bias. An introduction is provided at the very beginning. NUMBER will be replaced by the values stated in table 5, 6 and 7 .
Prompt Templates Beginning Instruction You are invited to participate in an experiment. Your task is to choose between option K and option U for each of the following lotteries. Here is lottery {i}: a) Used in Probability Mapping For option K: p% probability to win \$100. For option U: MARKER to win \$100. b) Used in PT Estimation For option U: NUMBER% probability to win \$NUMBER NUMBER% probability to win/lose \$NUMBER For option K: NUMBER% probability to win \$NUMBER NUMBER% probability to win/lose \$NUMBER End Instruction Please DO NOT REASON and DIRECTLY output your choice, by ONLY returning one of the following two labels: “[K]”, “[U]”. The answer is:
Figure 7: Templates for prompts used in the probability mapping and Prospect Theory estimation tasks. The design includes an initial instruction, task-specific lottery descriptions, and a fixed closing instruction to ensure direct model responses without reasoning.
Appendix C Detailed Experimental Results
In the main text, we presented selected key experimental results and visualizations. To provide a more comprehensive view of model performance across different rounds, we include in this appendix the full set of parameter estimates and model fit metrics.
Specifically, Tables 8, 9, and 10 report the estimates of parameters $\sigma$, $\lambda$, and $\gamma$ with their 95% confidence intervals for each model and round. Table 11 summarizes the models’ mean absolute errors (MAE) and McFadden $R^{2}$ values across rounds.
These additional data offer deeper insights into model behavior and the dynamics observed throughout the experiments.
Model $\sigma$ (95% CI)
baseline round1 round2 round3 round4
Llama-3.1-8B-Instruct 0.585(0.578,0.592)0.585\,(0.578,0.592) 0.605(0.599,0.612)0.605\,(0.599,0.612) 0.573(0.566,0.580)0.573\,(0.566,0.580) 0.463(0.455,0.473)0.463\,(0.455,0.473) 0.593(0.585,0.602)0.593\,(0.585,0.602)
Mistral-7B-Instruct-v0.3 0.534(0.526,0.543)0.534\,(0.526,0.543) 0.566(0.558,0.574)0.566\,(0.558,0.574) 0.571(0.564,0.580)0.571\,(0.564,0.580) 0.384(0.370,0.399)0.384\,(0.370,0.399) 0.444(0.431,0.454)0.444\,(0.431,0.454)
Qwen2.5-7B-Instruct 0.429(0.415,0.445)0.429\,(0.415,0.445) 0.010(0.010,0.074)0.010\,(0.010,0.074) 0.250(0.232,0.275)0.250\,(0.232,0.275) 0.010(0.010,0.071)0.010\,(0.010,0.071) 0.344(0.332,0.359)0.344\,(0.332,0.359)
Qwen2.5-14B-Instruct 0.503(0.495,0.511)0.503\,(0.495,0.511) 0.444(0.434,0.454)0.444\,(0.434,0.454) 0.409(0.398,0.421)0.409\,(0.398,0.421) 0.563(0.555,0.573)0.563\,(0.555,0.573) 0.589(0.581,0.600)0.589\,(0.581,0.600)
Qwen2.5-32B-Instruct 0.598(0.591,0.605)0.598\,(0.591,0.605) 0.569(0.561,0.576)0.569\,(0.561,0.576) 0.564(0.556,0.572)0.564\,(0.556,0.572) 0.647(0.640,0.655)0.647\,(0.640,0.655) 0.607(0.599,0.615)0.607\,(0.599,0.615)
Table 8: $\sigma$ estimates with 95% confidence intervals across different rounds for each model.
Model $\lambda$ (95% CI)
baseline round1 round2 round3 round4
Llama-3.1-8B-Instruct 0.010(0.010,0.125)0.010\,(0.010,0.125) 0.010(0.010,0.117)0.010\,(0.010,0.117) 0.010(0.010,0.130)0.010\,(0.010,0.130) 0.010(0.010,0.168)0.010\,(0.010,0.168) 0.010(0.010,0.112)0.010\,(0.010,0.112)
Mistral-7B-Instruct-v0.3 0.570(0.453,0.688)0.570\,(0.453,0.688) 0.010(0.010,0.132)0.010\,(0.010,0.132) 0.010(0.010,0.135)0.010\,(0.010,0.135) 2.260(2.060,2.445)2.260\,(2.060,2.445) 0.267(0.105,0.414)0.267\,(0.105,0.414)
Qwen2.5-7B-Instruct 0.010(0.010,0.584)0.010\,(0.010,0.584) 4.000(0.010,4.000)4.000\,(0.010,4.000) 4.000(3.526,4.000)4.000\,(3.526,4.000) 4.000(0.010,4.000)4.000\,(0.010,4.000) 4.000(3.736,4.000)4.000\,(3.736,4.000)
Qwen2.5-14B-Instruct 1.909(1.801,2.013)1.909\,(1.801,2.013) 1.919(1.784,2.070)1.919\,(1.784,2.070) 2.851(2.675,3.023)2.851\,(2.675,3.023) 2.191(2.094,2.295)2.191\,(2.094,2.295) 2.531(2.409,2.648)2.531\,(2.409,2.648)
Qwen2.5-32B-Instruct 1.213(1.133,1.295)1.213\,(1.133,1.295) 1.340(1.250,1.423)1.340\,(1.250,1.423) 0.953(0.866,1.036)0.953\,(0.866,1.036) 2.013(1.945,2.090)2.013\,(1.945,2.090) 1.815(1.736,1.905)1.815\,(1.736,1.905)
Table 9: $\lambda$ estimates with 95% confidence intervals across different rounds for each model.
Model $\gamma$ (95% CI)
baseline round1 round2 round3 round4
Llama-3.1-8B-Instruct 0.753(0.740,0.767)0.753\,(0.740,0.767) 0.750(0.737,0.762)0.750\,(0.737,0.762) 0.755(0.741,0.768)0.755\,(0.741,0.768) 1.020(0.994,1.053)1.020\,(0.994,1.053) 1.478(1.443,1.517)1.478\,(1.443,1.517)
Mistral-7B-Instruct-v0.3 0.577(0.565,0.590)0.577\,(0.565,0.590) 0.564(0.554,0.574)0.564\,(0.554,0.574) 0.535(0.527,0.545)0.535\,(0.527,0.545) 1.413(1.300,1.564)1.413\,(1.300,1.564) 1.332(1.266,1.414)1.332\,(1.266,1.414)
Qwen2.5-7B-Instruct 3.645(3.501,3.783)3.645\,(3.501,3.783) 1.156(0.010,4.000)1.156\,(0.010,4.000) 0.985(0.870,1.174)0.985\,(0.870,1.174) 1.035(0.010,3.498)1.035\,(0.010,3.498) 0.987(0.926,1.053)0.987\,(0.926,1.053)
Qwen2.5-14B-Instruct 0.896(0.875,0.919)0.896\,(0.875,0.919) 0.933(0.905,0.967)0.933\,(0.905,0.967) 0.983(0.944,1.023)0.983\,(0.944,1.023) 1.455(1.414,1.503)1.455\,(1.414,1.503) 1.742(1.693,1.803)1.742\,(1.693,1.803)
Qwen2.5-32B-Instruct 0.867(0.851,0.884)0.867\,(0.851,0.884) 0.778(0.762,0.793)0.778\,(0.762,0.793) 0.812(0.797,0.828)0.812\,(0.797,0.828) 0.614(0.605,0.623)0.614\,(0.605,0.623) 0.664(0.653,0.674)0.664\,(0.653,0.674)
Table 10: $\gamma$ estimates with 95% confidence intervals across different rounds for each model.
Model      MAE     McFadden $R^{2}$
     baseline round1 round2 round3 round4     baseline round1 round2 round3 round4
Llama-3.1-8B-Instruct      0.332 0.320 0.326 0.242 0.155     0.092 0.121 0.090 0.082 0.335
Mistral-7B-Instruct-v0.3      0.155 0.202 0.196 0.067 0.075     0.132 0.162 0.190 0.056 0.130
Qwen2.5-7B-Instruct      0.047 0.125 0.159 0.387 0.157     0.116 0.000 0.011 -0.001 0.032
Qwen2.5-14B-Instruct      0.257 0.152 0.069 0.150 0.157     0.067 0.070 0.075 0.227 0.336
Qwen2.5-32B-Instruct      0.161 0.127 0.129 0.224 0.256     0.225 0.195 0.211 0.139 0.088
Table 11: Mean absolute error (MAE) and McFadden $R^{2}$ across different rounds for each model.
Appendix D Discussion and Implications
Our findings reveal challenging difficulties in applying human-centric cognitive frameworks, especially Prospect Theory (PT), to LLM decision-making. Different models display distinct interpretations of epistemic uncertainty markers, leading to divergent decision behaviors. Introducing these markers into the decision-making framework substantially alters LLM choices.
Our results suggest that LLMs may not inherently understand risk in human-like ways; their responses often reflect statistical training artifacts rather than cognitively grounded reasoning. We recommend conducting regression analyses and goodness-of-fit tests before applying human cognitive models to LLMs.
In real-world applications (e.g., medical diagnosis or financial advice), LLMs may give inconsistent recommendations when probabilistic language varies, posing reliability concerns. We recommend establishing consistent standards for expressing uncertainty in LLM-driven decision systems.
Furthermore, larger LLMs tend to exhibit more PT-like decision behavior, with PT parameters more closely aligned to human estimates. We recommend using LLMs with at least 14B parameters when integrating PT into decision-making systems.
Appendix E Failure Cases Analysis
In our initial implementation of the marker mapping experiment, we adopted the set of epistemic markers from Table 6 (“Human Judgements of Templates Based on Reliability”) in Zhou et al. (2024). These markers were originally designed to test both human and LMs’ judgments of the reliability conveyed by these expressions. However, when directly applied to our economic decision-making setting, the resulting mappings for LLMs were unexpectedly unstable and, in some cases, counterintuitive.
Failure Case of Markers Mapping user: Here is lottery 2: For option K: 15% probability to win \$100. For option U: It’s extremely certain to win \$100. Please DO NOT REASON and DIRECTLY output your choice,
by ONLY returning one of the following two labels: “[K]”, “[U]”.
The answer is:
  assistant: K
Figure 8: An example of mapping high-certainty markers to low probablity. In this case, Qwen2.5-7B-Instruct preferred option K when $p=15\%$ and option U was described using “It’s extremely certain to”, implying a far lower internal probability than expected.
We summarize two major issues observed in the experimental outcomes:
(1) Highly oscillatory choice patterns. Ideally, the number of times the model selects option K should increase monotonically with $p$, yielding a single and well-defined switching point. In practice, the selection curves were often non-monotonic, with multiple apparent switching points, which made the mapping probability $p_{\text{mapping}}$ ill-defined. An example is shown in Figure 9.
Figure 9: An example of non-monotonic result with multiple switching points. This result comes from marker “It’s undoubtedly to” with Qwen2.5-7B-Instruct model over 256 trials. The blue line shows counts of option K selections, and the yellow dashed line marks half the trials. The crosses denote switching points where choices for K and U are equal.
(2) Severe semantic mismatches for high-certainty markers. Some markers, such as “It’s extremely certain to”, convey very high certainty in human interpretation, but were mapped by the model to surprisingly low numerical probabilities. An example is shown in Figure 8.
We hypothesize that these issues may stem from several factors:
(1) Marker length and syntactic complexity. The selected markers were often not single words or short phrases, but full clausal structures. This may introduce additional semantic and syntactic cues unrelated to uncertainty, thereby interfering with probability interpretation.
(2) Shift from first-person to third-person framing. The original markers in Zhou et al. (2024) were presented in the first person (e.g., “I am not confident, maybe it’s…”), whereas our experiment reformulated them into third-person expressions (e.g., “It’s not confident, maybe can…”). This change in perspective may alter the perceived epistemic force of the marker for the model.
(3) Intrinsic instability of epistemic markers. Even for human interpretation, such markers are context-dependent and inherently imprecise Liu et al. (2025b); Guo et al. (2025); Wang et al. (2025); Zheng et al. (2025a, b). Their probability mapping by LLMs in economic decision-making contexts may exhibit fundamental reliability flaws Zheng et al. (2025c).
These limitations motivated the redesign of our marker set and prompt formulation in subsequent experiments, aiming for shorter, syntactically simpler, and context-consistent expressions.